# Targeted Data Fusion for Causal Survival Analysis Under Distribution Shift

### v1

@@ 2.3.1 Local Sites
For a given $k\in\{0,1,\dots,K-1\}$, if there exists an $\eta\in(0,\infty)$ such that $\min\{\pi^{k}(a\mid\mathbf{x}),G^{k}(t\mid a,\mathbf{x})\}\geq 1/\eta$, $\omega^{k,0}(\mathbf{x})\leq\eta$, and $S^{k}(t\mid a,\mathbf{x})>0$ for $\mathbb{P}$-almost every $\mathbf{x}$, then $\theta^{0}(t,a)$ is a pathwise differentiable parameter given $t\in[0,\tau]$ and $a\in\{0,1\}$, in a nonparametric model with semiparametric EIF $\varphi^{*k,0}_{t,a}(\mathcal{O};S^{k},S^{0},G^{k},\pi^{k},\omega^{k,0})=\varphi^{*k,0}_{t,a}(\mathcal{O};\mathbb{P})=\varphi^{k,0}_{t,a}(\mathcal{O};\mathbb{P})-\theta^{0}(t,a)$, where $\varphi^{k,0}_{t,a}(\mathcal{O};\mathbb{P})=$
$\frac{\mathbb{I}(R=0)}{\mathbb{P}(R=0)}S^{0}(t\mid a,\mathbf{X})-\frac{\mathbb{I}(R=k)}{\mathbb{P}(R=k)}\omega^{k,0}(\mathbf{X})S^{k}(t\mid a,\mathbf{X})$
×$\quad\times\frac{\mathbb{I}(A=a)}{\pi^{k}(a\mid\mathbf{X})}\bigg{\{}\frac{\mathbb{I}(Y\leq t,\Delta=1)}{S^{k}(Y\mid a,\mathbf{X})G^{k}(Y\mid a,\mathbf{X})}$
$\quad\quad-\int_{0}^{t\wedge Y}\frac{\Lambda^{k}(du\mid a,\mathbf{X})}{S^{k}(u\mid a,\mathbf{X})G^{k}(u\mid a,\mathbf{X})}\bigg{\}},$
@@
A site-specific estimator can be motivated from the EIF above, namely $\widehat{\theta}^{k,0}_{n}(t,a)=\mathbb{P}_{n}[\widehat{\varphi}^{k,0}{t,a}(\mathcal{O};\widehat{\mathbb{P}})]$. Of note regarding $\widehat{\theta}^{k,0}_{n}(t,a)$: (i) the “anchor” term in the EIF uses target-site data ($\mathbb{I}(R=0)$), while the “augmentation” term leverages site-$k$ data ($\mathbb{I}(R=k)$); (ii) all information is obtained independently from each site except for the density ratio $\omega^{k,0}(\mathbf{X})$, which is estimated by sharing only coarse summary statistics under flexible models (Han et al., 2021), thereby adhering to data sharing constraints; and (iii) for $S^{k}$ in the augmentation term, we train an $S^{0}$ model on the target site and apply its predictions to site $k$, since $S^{0}$ and $S^{k}$ are exchangeable under partial CCOD. Crucially, if partial CCOD is violated, we can detect site heterogeneity by the difference between $\widehat{\theta}^{k,0}_{n}(t,a)$ and $\widehat{\theta}^{0}_{n}(t,a)$.
@@
To ensure the monotonicity of the estimated site-specific survival curves, we invoke isotonic regression techniques (Westling et al., 2020), which enforce a non-increasing constraint on the survival estimates, thereby maintaining their logical consistency over time.
@@ 2.3.2 Aggregation across sites
To aggregate information from the target and source sites, for a fixed time $t$ and treatment $a$, we use an $L_{1}$-penalization procedure to calculate federated weights for each site. We define $\widehat{\chi}_{n,t,a}^{k,0}=\widehat{\theta}^{k,0}(t,a)-\widehat{\theta}^{k}(t,a)$, and write $\boldsymbol{\eta}_{t,a}=(\eta^{0}_{t,a},\eta^{1}_{t,a},\dots,\eta^{K-1}_{t,a})$ for the vector of federated weights. We estimate $\boldsymbol{\eta}_{t,a}$ data-adaptively by minimizing the following $L_{1}$-penalized objective function:
$Q(\boldsymbol{\eta}_{t,a})$ $=\mathbb{P}_{n}\left[\left\{\widehat{\varphi}^{*0}_{t,a}(\mathcal{O};\widehat{\mathbb{P}})-\sum_{k=1}^{K-1}\eta^{k}_{t,a}\widehat{\varphi}^{*k,0}_{t,a}(\mathcal{O};\widehat{\mathbb{P}})\right\}^{2}\right]$
$\qquad+\frac{1}{n}\lambda\sum_{k=1}^{K-1}|\eta^{k}_{t,a}|(\widehat{\chi}_{n,t,a}^{k,0})^{2},$ (1)
subject to $0\leq\eta^{k}_{t,a}\leq 1$, where $\widehat{\chi}_{n,t,a}^{k,0}=\widehat{\theta}^{k,0}(t,a)-\widehat{\theta}^{k}(t,a)$, for all $k=0,1,\dots,K-1$ and $\sum_{k=0}^{K-1}\eta^{k}_{t,a}=1$, and $\lambda$ is a tuning parameter that controls the bias-variance trade-off and is chosen by cross-validation.
The objective function follows the principle that one should ascribe higher (resp. lower) weight to the more (resp. less) informative sites. Heuristically, our approach anchors on the nonparametric estimate $\widehat{\theta}^{0}_{n}(t,a)$ and weights site $k$ when it is deemed similar enough to the target site. Then, we compute $\widehat{\theta}^{\text{fed}}_{n}(t,a)$, our “federated” (FED) estimator as the weighted average of the site-specific survival estimates:
@@
We present the details for implementing the FED estimator in Algorithm 1. Figure 1 further gives an intuitive visualization of the proposed federated algorithm.
Figure 1: Illustration of the Federated Algorithm: Each site has its underlying survival functions, which are first estimated site-specifically. The algorithm is then applied over all $a\in\{0,1\}$ and $t\in[0,\tau]$ (using a finite grid for the time interval in practice) to derive treatment- and time-specific federated weights for the target and source sites, ultimately producing the weighted average survival functions (the purple step).
@@ 2.3.3 Theoretical results of the federated estimator
Finally, we establish the pointwise RAL property and the efficiency gain of the federated estimator in Theorem 2.12.
@@
If the regularity conditions (D.1, D.2 and D.4) specified in Appendix D.2 hold, the federated estimator $\widehat{\theta}^{\text{fed}}_{n}(t,a)$ is pointwise RAL given $t\in[0,\tau]$ and $a\in\{0,1\}$,
$\sqrt{n/\widehat{\mathcal{V}}}\left\{\widehat{\theta}^{\text{fed}}_{n}(t,a)-\theta^{0}(t,a)\right\}\to_{d}\mathcal{N}(0,1),$
with the variance consistently estimated as $\widehat{\mathcal{V}}$ (see Appendix D.3). The variance of $\widehat{\theta}^{\text{fed}}_{n}(t,a)$ is no larger than that of the estimator based on the target data only, $\widehat{\theta}^{0}_{n}(t,a)$. Further, if there exists some source sites that consistently estimate $\theta^{0}(t,a)$, and the oracle selection conditions specified in Appendix D.3 are satisfied, the variance of $\widehat{\theta}^{\text{fed}}_{n}(t,a)$ is strictly smaller than $\widehat{\theta}^{0}_{n}(t,a)$.
Algorithm 1 Federated Estimator for Multi-Source Causal Survival Analysis.
1:  Input: Observed multi-source right-censored data $\mathcal{O}=\{(\mathbf{X}_{i},A_{i},Y_{i},\Delta_{i},R_{i}),i=1,\dots,n\}$, a time horizon $\tau>0$, and a time grid divided by an $\epsilon>0$, i.e., $\{0,\epsilon,2\epsilon,\dots,\tau\}$ is a finite and fine grid for interval $[0,\tau]$.
2:  Output: Estimated treatment-specific curves $\widehat{\theta}^{\text{fed}}_{n}(t,a)$ and variance $\widehat{\mathcal{V}}=\widehat{\mathcal{V}}(\widehat{\theta}^{\text{fed}}_{n}(t,a))$ for $a\in\{0,1\}$ and $t\in[0,\tau]$.
3:  for $(t,a)\in\{0,\epsilon,2\epsilon,\dots,\tau\}\times\{0,1\}$ do
8:  end for
9:  Return:
$\widehat{\theta}^{\text{fed}}_{n}(t,a)=\sum_{k=0}^{K-1}\widehat{\eta}^{k}_{t,a}\widehat{\theta}^{k,0}_{n}(t,a),\text{ and }\widehat{\mathcal{V}}=\mathcal{V}(\widehat{\boldsymbol{\eta}}_{t,a})^{\dagger}\text{ for all }(t,a)\in\{0,\epsilon,2\epsilon,\dots,\tau\}\times\{0,1\}.$
$\dagger$: the formula of $\widehat{\mathcal{V}}$ is specified in (D.3) in Appendix D.3 by replacing the fixed ${\boldsymbol{\eta}}_{t,a}$ with the estimated value $\widehat{\boldsymbol{\eta}}_{t,a}$ here.
@@ 3 Numerical Experiments
We conduct extensive simulations to evaluate our “CCOD” and “FED” estimators against competing methods: target-site-only (TGT), simple pooling (POOL), and inverse variance weighting (IVW). TGT uses only target site data ($R=0$), POOL pools data from all sites without adjustment, and IVW computes a weighted average of site-specific estimators, with weights based on inverse variance. We also consider “FED (BOOT),” as another version of FED, which applies bootstrap (200 replicates) when deriving the federated weights using (2.3.2), aiming to reduce uncertainty associated with the estimated influence functions.
3.1 Simulation set-up
We conduct $M=500$ Monte Carlo replications, with $n=\sum_{k=1}^{K}n_{k}$ observations distributed across $K=5$ sites. Each site $k\in\{0,1,\dots,4\}$ has $n_{k}$ observations, with site $k=0$ as the target site. We fix the target site sample size at $n_{0}=300$ and vary $n_{k}\in\{300,600,1000\}$ for all $k=1,\dots,4$, corresponding to limited, moderate, and abundant source data. Covariate, treatment, and outcome generation details are provided in Appendix A.1. The true survival curves are derived by averaging over a super-population of size $n_{\text{super}}=10^{8}$ under the target site’s data generation.
@@ 3.1 Simulation set-up
Figure 2: Simulation results evaluated at time $=$ day 90, under $n_{0}=300$ and $n_{k}=600$ ($k\geq 1$).
@@ 3.2 Results
For succinctness, we present the simulation results only under day 90 and $n_{k}=600$ ($k\geq 1$) in this section. The full simulation results can be found in Appendix A.2. The main results (Figure 2) show that FED and FED (BOOT) consistently exhibit biases centered around zero (panel (A)). Across all site heterogeneity and treatment scenarios, TGT proves the least efficient, while FED and FED (BOOT) achieve a 5%–16% lower RMSE by the RRMSE in panel (B), confirming that our federated methods are consistent with improved efficiency. Furthermore, both federated methods provide 95% CIs that are no wider than TGT’s and produce CPvalues near the nominal 95% level. In addition, Figure 5 in Appendix A.2 displays the federated weights across different scenarios. The figure includes a panel showing average weights over time across 500 simulations, as well as a panel with scatterplots of weights versus $(\widehat{\chi}_{n,t,a}^{k})^{2}$. These plots reveal that federated weights generally decrease as $(\widehat{\chi}_{n,t,a}^{k})^{2}$ increases. The target site often receives higher weights when covariate and outcome shifts are present. These weights can also vary at different time points based on the similarity of survival function values at a given time.
In contrast, although POOL and IVW yield smaller variance (as shown by narrower boxplots and CIs), they are highly biased under Covariate Shift, Outcome Shift, and All Shift scenarios, resulting in elevated RRMSE and severely low CP%. However, these methods remain robust under Censoring Shift, consistent with the notion that censoring is estimated separately on each site and plays an ancillary role.
Lastly, the CCOD estimator stands out for its robustness and efficiency under Covariate Shift, aligning with its theoretical strengths (consistent, semiparametrically efficient, and accommodating covariate shift). Nonetheless, it can fail under Outcome or All Shifts, because its validity depends on the CCOD assumption.
@@ 4 Real Data Analysis
Figure 3: Panel (A): Age, standardized risk score, and weight are summarized by mean (standard error). Panel (B): Thinner, lighter lines depict estimated time-specific federated weights, while bolder lines show smoothed weights using locally weighted regression (Cleveland & Devlin, 1988) to illustrate trend patterns over the study period. Panel (C): Missing values at early time points in IVW indicate large IVW weights. Relative efficiency, defined as the ratio of the estimated standard deviation to that of the TGT estimator, is presented for three specific days: 148, 330, and 512.
@@
We apply our methods to two coordinated randomized trials, HVTN 704/HPTN 085 and HVTN 703/HPTN 081 (Corey et al., 2021; Ning et al., 2023), which together enrolled 4,611 participants to evaluate whether a broadly neutralizing monoclonal antibody (bnAb) reduces the risk of HIV-1 acquisition. The endpoint was an HIV diagnosis by week 80–a rare event with a 3.77% incidence. The overall rate of loss to follow-up in this study is low (less than 10% in each treatment group), as indicated by Corey et al. (2021).
@@
We divide the data into four subsets by region: (i) SA: South Africa, (ii) OA: other sub-Saharan African countries, (iii) BP: Brazil or Peru, and (iv) US: the United States or Switzerland. Participants in (i) and (ii) were women, whereas those in (iii) and (iv) were cisgender men and transgender individuals, potentially introducing distribution shifts across regions. Because HIV diagnoses were infrequent, we only use a 2-fold cross-fitting, and estimate conditional survival for both event and censoring processes by an ensemble of Kaplan–Meier, Cox regression and survival random forest models via the survSuperLearner package (Westling et al., 2023). The propensity score and density ratio (used in federated method) models are fitted by the ensemble of logistic regression and LASSO from super learner (van der Laan et al., 2007). In all these nuisance models, age (year), a standardized risk score built by machine learning, and weight (kg) at baseline are included as covariates.
@@
In this section, we focus on the SA region as our target site. We include the results of treating other regions as target site (Figure 10) as well as a direct comparison of site-specific curves (Figure 9) in Appendix B. Panel (A) of Figure 3 summarizes baseline covariates and the event by region and treatment group. OA’s summary statistics are similar SA’s, whereas BP and US differ notably in mean standardized risk scores, weight, and HIV prevalence, indicating covariate and possibly outcome shifts compared to SA. This aligns with panel (B) for federated weights over study time, where SA receives the highest weights overall, followed by OA, US, and BP.
@@
Panel (C) shows that the TGT, FED, and FED (BOOT) methods produce similar survival curves. However, TGT sometimes has wider confidence intervals due to lower efficiency and lacks interval estimates at certain early time points when variance estimates are unavailable or too small. In contrast, FED and FED (BOOT) can provide some interval estimates in these cases. The CCOD method generates comparable survival curves with narrower confidence intervals and smaller variances. The efficiency gains observed with FED, FED (BOOT), and CCOD are consistent with our simulation results in Section 3. On the other hand, IVW and POOL methods, while offering more efficient estimates as indicated by lower relative efficiency values, diverge from the other methods from the trends on their survival curves, suggesting potential bias due to covariate and outcome distribution shifts when targeting the South Africa population.

### v2

@@ 2.2 Efficient estimation under CCOD and data sharing
Remark 2.9.
Among the global nuisance functions $\bar{S}$, $\bar{\pi}$, and $\bar{G}$, only $\bar{S}$ is assumed (based on CCOD) to equal each corresponding site-specific function (in this case, $S^{k}$, for all $k$). The inclusion of $\bar{\pi}$ and $\bar{G}$ in Theorem 2.7 arises purely from the EIF derivation and does not require homogeneity assumptions on the distributions of $A$ or $C^{(a)}$ across sites. In practice, $A$ and $C^{(a)}$ are modeled as functions of $\mathbf{X}$ using the pooled sample. Moreover, inspecting the forms of the EIFs in Proposition 2.4 and Theorem 2.7, it is clear that the CCOD estimator, unlike the target-only estimator, incorporates data from all sites: while the “anchor” term of $\widehat{\theta}^{\text{CCOD}}_{n}(t,a)$ relies on target data only, the “augmentation” term incorporates information from all sites.
@@ 2.3.1 Local Sites
For $k\in\{0,1,\dots,K-1\}$, $\theta^{0}(t,a)$ is a pathwise differentiable parameter given $t\in[0,\tau]$ and $a\in\{0,1\}$, with semiparametric EIF—under the working assumption that $S^{k}\equiv S^{0}$— given by $\varphi^{*k,0}_{t,a}(\mathcal{O};S^{k},S^{0},G^{k},\pi^{k},\omega^{k,0})=\varphi^{*k,0}_{t,a}(\mathcal{O};\mathbb{P})=\varphi^{k,0}_{t,a}(\mathcal{O};\mathbb{P})-\theta^{0}(t,a)$, where
$\varphi^{k,0}_{t,a}(\mathcal{O};\mathbb{P})=\frac{\mathbb{I}(R=0)}{\mathbb{P}(R=0)}S^{0}(t\mid a,\mathbf{X})-\frac{\mathbb{I}(R=k)}{\mathbb{P}(R=k)}\omega^{k,0}(\mathbf{X})S^{k}(t\mid a,\mathbf{X})\frac{\mathbb{I}(A=a)}{\pi^{k}(a\mid\mathbf{X})}\mathcal{H}_{t,a}(\mathcal{O};S^{k},G^{k}),$
@@
A site-specific estimator can be motivated from the EIF above, namely $\widehat{\theta}^{k,0}_{n}(t,a)=\mathbb{P}_{n}[\widehat{\varphi}^{k,0}_{t,a}(\mathcal{O};\widehat{\mathbb{P}})]$. Of note regarding $\widehat{\theta}^{k,0}_{n}(t,a)$: (i) the “anchor” term in the EIF uses target-site data ($\mathbb{I}(R=0)$), while the “augmentation” term leverages site-$k$ data ($\mathbb{I}(R=k)$); (ii) all information is obtained independently from each site except for the density ratio $\omega^{k,0}(\mathbf{X})$, which is estimated by sharing only coarse summary statistics under flexible models [25], thereby adhering to data sharing constraints; and (iii) for $S^{k}$ in the augmentation term, we train an $S^{0}$ model on the target site and apply its predictions to site $k$, since $S^{0}$ and $S^{k}$ are exchangeable under partial CCOD. Crucially, if partial CCOD is violated, we can detect site heterogeneity by the difference between $\widehat{\theta}^{k,0}_{n}(t,a)$ and $\widehat{\theta}^{0}_{n}(t,a)$.
@@
To ensure the monotonicity of the estimated site-specific survival curves, we invoke isotonic regression techniques [49], which enforce a non-increasing constraint on the survival estimates, thereby maintaining their logical consistency over time.
@@ 2.3.2 Aggregation across sites
To aggregate information from the target and source sites, for a fixed time $t$ and treatment $a$, we use an $L_{1}$-penalization procedure to calculate federated weights for each site. We define $\widehat{\chi}_{n,t,a}^{k,0}=\widehat{\theta}^{k,0}(t,a)-\widehat{\theta}^{0}(t,a)$, and write $\boldsymbol{\eta}_{t,a}=(\eta^{0}_{t,a},\eta^{1}_{t,a},\dots,\eta^{K-1}_{t,a})$ for the vector of federated weights. We estimate $\boldsymbol{\eta}_{t,a}$ data-adaptively by minimizing the following $L_{1}$-penalized objective function:
$Q(\boldsymbol{\eta}_{t,a})=\mathbb{P}_{n}\left[\left\{\widehat{\varphi}^{*0}_{t,a}(\mathcal{O};\widehat{\mathbb{P}})-\sum_{k=1}^{K-1}\eta^{k}_{t,a}\widehat{\varphi}^{*k,0}_{t,a}(\mathcal{O};\widehat{\mathbb{P}})\right\}^{2}\right]+\frac{1}{n}\lambda\sum_{k=1}^{K-1}|\eta^{k}_{t,a}|(\widehat{\chi}_{n,t,a}^{k,0})^{2},$ (2)
subject to $0\leq\eta^{k}_{t,a}\leq 1$, where $\widehat{\chi}_{n,t,a}^{k,0}=\widehat{\theta}^{k,0}(t,a)-\widehat{\theta}^{0}(t,a)$, for all $k=0,1,\dots,K-1$ and $\sum_{k=0}^{K-1}\eta^{k}_{t,a}=1$, and $\lambda$ is a tuning parameter that controls the bias-variance trade-off and is chosen by cross-validation.
The objective function follows the principle of assigning higher (resp. lower) weight to more (resp. less) informative sites. Heuristically, our approach anchors on the nonparametric estimate $\widehat{\theta}^{0}_{n}(t,a)$ and weights site $k$ when it is deemed similar to the target site. We then compute $\widehat{\theta}^{\text{fed}}_{n}(t,a)$, our “federated” (FED) estimator, as the weighted average of the site-specific survival estimates.
@@
Figure 2 below gives an intuitive visualization of the proposed FED algorithm. We also present all details for the FED method in Algorithm 1 in Appendix C.
Figure 2: Illustration of the Federated Algorithm. Each site has its underlying survival functions, which are first estimated in a site-specific manner. Then, for all $a\in\{0,1\}$ and $t\in[0,\tau]$, treatment- and time-specific federated weights for the target and source sites are derived, ultimately producing the weighted average survival functions (the purple step).
@@ 2.3.3 Theoretical results of the federated estimator
We establish the asymptotic properties and efficiency gain of the federated estimator in the following result.
@@
If Conditions D.6D.8 in Appendix D.2 (i.e., consistency, positivity, and certain rate conditions on the product of prediction errors of the nuisance functions at each local site) and oracle selection conditions for the federated weights $\boldsymbol{\eta}_{t,a}$ in Appendix D.3 hold, then the federated estimator $\widehat{\theta}^{\text{fed}}_{n}(t,a)$ at each $(t,a)\in[0,\tau]\times\{0,1\}$ has asymptotic distribution
$\sqrt{n/\widehat{\mathcal{V}}_{t,a}}\left\{\widehat{\theta}^{\text{fed}}_{n}(t,a)-\theta^{0}(t,a)\right\}\to_{d}\mathcal{N}(0,1),$
where $\widehat{\mathcal{V}}_{t,a}$ is a consistent estimator for the asymptotic variance of $\widehat{\theta}^{\text{fed}}_{n}(t,a)$ (see Appendix D.3 for its form), which is no greater than the asymptotic variance of the target-only estimator, $\widehat{\theta}^{0}_{n}(t,a)$. Moreover, if any source site provides a consistent estimate of $\theta^{0}(t,a)$, then the asymptotic variance of $\widehat{\theta}^{\text{fed}}_{n}(t,a)$ is strictly smaller than that of $\widehat{\theta}^{0}_{n}(t,a)$.
All technical proofs are presented in Appendices D.2 and D.3. Here, we briefly outline the key theoretical insights and their justifications. To quantify the efficiency gain afforded by our federated method, we adopt an approach analogous to that in existing federated learning studies [25, 33, 27].
Letting $\mathcal{S}=\{1,\dots,K-1\}$ denote the set of source sites, we define the oracle selection space for $\boldsymbol{\eta}_{t,a}$ and the corresponding weight space as, respectively, $\mathcal{S}^{*}_{t,a}=\{k\in\mathcal{S}:\theta^{k}(t,a)=\theta^{0}(t,a)\}$ and $\mathbb{R}^{S^{*}_{t,a}}=\{\boldsymbol{\eta}_{t,a}\in\mathbb{R}^{K-1}:\eta^{j}_{t,a}=0,\forall j\not\in\mathcal{S}^{*}_{t,a}\}$. Appendix D.3 delineates conditions under which our federated approach can recover the following optimal weights by solving (2):
$\bar{\boldsymbol{\eta}}_{t,a}=\underset{\eta^{k}_{t,a}=0,\forall k\not\in\mathcal{S}^{*}_{t,a}}{\text{arg min}}\mathcal{V}_{t,a}(\boldsymbol{\eta}_{t,a}),$
where $\mathcal{V}_{t,a}(\boldsymbol{\eta}_{t,a})$ is the asymptotic variance of the federated estimator under a fixed weight vector $\boldsymbol{\eta}_{t,a}$. Since the target-only estimator corresponds to a special case $\boldsymbol{\eta}_{t,a}=(1,0,\dots,0)$ (assigning weight 1 only to the target site), its asymptotic variance is no larger than that of any federated estimator. Moreover, the key rationale behind the federated weighting is that if the bias term $\widehat{\chi}_{n,t,a}^{k,0}$ remains asymptotically non-zero, the corresponding weight $\eta^{k}_{t,a}$ converges to zero.
@@ 3 Numerical Experiments
We conduct simulations to evaluate our “CCOD” and “FED” estimators against competing methods: target-site-only (TGT), simple pooling (POOL), and inverse variance weighting (IVW). TGT uses only target site data ($R=0$), POOL pools data from all sites without adjustment, and IVW computes a weighted average of site-specific estimators, with weights based on inverse variance. We also consider “FED (BOOT),” as another version of FED, which applies bootstrap (200 replicates) when deriving the federated weights using (2), aiming to reduce uncertainty associated with the estimated influence functions.
3.1 Data generation and performance metrics
We conduct $M=500$ Monte Carlo replications, with $n=\sum_{k=1}^{K}n_{k}$ observations distributed across $K=5$ sites. Each site $k\in\{0,1,\dots,4\}$ has $n_{k}$ observations, with site $k=0$ as the target site. We fix the target site sample size at $n_{0}=300$ and vary $n_{k}\in\{300,600,1000\}$ for all $k=1,\dots,4$, corresponding to limited, moderate, and abundant source data. Covariate, treatment, and outcome generation details are provided in Appendix A.1. The true survival curves are derived by averaging over a simulated population of size $n_{\text{super}}=10^{8}$ generated from the target site distribution.
@@ 3.2 Results
For succinctness, we present the simulation results only under day 90 and $n_{k}=600$ ($k\geq 1$) in this section. The full simulation results can be found in Appendix A.2. The main results (Figure 3) show that FED and FED (BOOT) consistently exhibit biases centered around zero (panel (A)). Across all site heterogeneity and treatment scenarios, TGT proves the least efficient, while FED and FED (BOOT) achieve a 5%–16% lower RMSE by the RRMSE in panel (B), confirming that our federated methods are consistent with improved efficiency. Furthermore, both federated methods provide 95% CIs that are no wider than TGT’s and produce CPvalues near the nominal 95% level. In addition, Figure 6 in Appendix A.2 displays the federated weights across different scenarios. The figure includes a panel showing average weights over time across 500 simulations, as well as a panel with scatterplots of weights versus $(\widehat{\chi}_{n,t,a}^{k,0})^{2}$. These plots reveal that federated weights generally decrease as $(\widehat{\chi}_{n,t,a}^{k,0})^{2}$ increases. The target site often receives higher weights when covariate and outcome shifts are present. These weights can also vary at different time points based on the similarity of survival function values at a given time.
Figure 3: Simulation results evaluated at day 90, under $n_{0}=300$ and $n_{k}=600$, $k=1,2,3,4$.
In contrast, although POOL and IVW yield smaller variance (as shown by narrower boxplots and CIs), they are highly biased under Covariate, Outcome, and All Shifts, resulting in elevated RRMSE and severely low CP%. However, these methods remain robust under Censoring Shift, consistent with the notion that censoring is estimated separately on each site and plays an ancillary role.
Lastly, the CCOD estimator stands out for its consistency and efficiency under Covariate Shift, aligning with its theoretical strengths (consistent, semiparametrically efficient, and accommodating covariate shift), but it can fail under Outcome or All Shifts due to it reliance on the CCOD assumption.
@@ 4 Real Data Analysis
We applied our methods to two coordinated randomized trials, HVTN 704/HPTN 085 and HVTN 703/HPTN 081 [12, 39], which together enrolled 4,611 participants to evaluate whether a broadly neutralizing monoclonal antibody (bnAb) reduces the risk of HIV-1 acquisition. The endpoint was an HIV diagnosis by week 80–a rare event with a 3.77% incidence. The overall rate of loss to follow-up in this study is low (less than 10% in each treatment group), as indicated by Corey et al. [12].
@@
We divided the data into four subsets by region: (i) SA: South Africa, (ii) OA: other sub-Saharan African countries, (iii) BP: Brazil or Peru, and (iv) US: the United States or Switzerland. Participants in (i) and (ii) were women, whereas those in (iii) and (iv) were cisgender men and transgender individuals, potentially introducing important distribution shifts across regions. Because HIV diagnoses were infrequent, we only used 2-fold cross-fitting, and estimated conditional survival for both event and censoring processes by an ensemble of Kaplan–Meier, Cox regression and survival random forest models via the survSuperLearner package [48]. The propensity score and density ratio (used in federated method) models were fitted by the ensemble of logistic regression and LASSO using the Super Learner [45]. In all these nuisance models, age (year), a standardized risk score built by machine learning, and weight (kg) at baseline were included as covariates.
Figure 4: Data analysis results (South Africa as the target site). Panel (A): Missing values early in IVW indicate large IVW weights. Relative efficiency, defined as the ratio of the estimated standard deviation to that of the TGT estimator, presented for three time points: 148, 330, and 512 days. Panel (B): Time-specific federated weights and smoothed weights using locally weighted regression [10].
@@
In this section, we focus on the SA region as our target site (Figure 4). Additional results—including analyses treating other regions as the target (Figure 11), direct comparisons of site-specific survival curves (Figure 10), and summaries of baseline covariates and event rates by region and treatment group (Table 1)—are provided in Appendix B. Table 1 shows that OA’s summary statistics closely mirror those of SA, whereas BP and US differ markedly in mean standardized risk scores, weight, and HIV prevalence, indicating covariate—and possibly outcome—shifts relative to SA. This pattern aligns with panel (B), where SA overall receives the highest federated weights over time, followed by OA, US, and BP.
@@
Panel (A) illustrates that the TGT, FED, and FED (BOOT) methods yield nearly identical survival curves. However, TGT occasionally has wider confidence intervals—reflecting lower efficiency—and fails to produce interval estimates at some early time points when its variance estimates are unavailable or too small. In contrast, FED and FED (BOOT) still provide interval estimates in those cases. The CCOD method generates comparable curves with narrower confidence intervals and lower variance. These efficiency gains for FED, FED (BOOT), and CCOD align with our simulation results in Section 3. By comparison, the IVW and POOL methods—while exhibiting higher nominal efficiency (lower relative-efficiency values)—diverge from the other methods’ survival trends, suggesting potential bias arising from covariate and outcome distribution shifts when targeting the SA population.